Find the number of ways of selecting a cricket team of 11 players from 7 batsmen and 6
bowlers such that there will be at least 5 bowlers in the team.

Grade:upto college level

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Given there should be atleast 5 bowlers in the team
so possible ways are 5 bowlers,6 batsmen and 6 bowlers 5 batsmen
=7c6*6c5+7c5*6c6=42+21=63
Thanks and Regards
Shaik Aasif
askIITians faculty